# Bandit and Delayed Feedback in Online Structured Prediction

### v1

@@ G.1 Setup
As the algorithm $\mathsf{ALG}$ to update the linear estimator, we employ the adaptive OGD in Section 3.2. Following Van der Hoeven et al. (2021), we use the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\left(10^{-8}+\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}\right)}}$ and no orthogonal projection in conducted in OGD. Since $B=\mathrm{diam}(\mathcal{W})$ is unknown, we fixed $B=10$ for the experiments, regardless of whether this value represents the actual diameter. All other parameters were set according to theoretical values. Under these parameter settings, we repeated experiments for $20$ times.
G.2 Results
The results on synthetic data are provided in Figure 2. Our algorithm achieves comparable or better results than the comparison methods for datasets when $K\leq 24$. By contrast, when $K=48,96$, the cumulative losses of our algorithm are larger compared to the other two algorithms, Gaptron with the hinge loss and Gappletron with the logistic loss. Note that these experimental results are consistent with existing theoretical results since for larger values of $K$ the upper bound on the cumulative loss of Gappletron is better than our algorithm due to differences in the surrogate loss function (see the discussion in Appendix C for details). Nevertheless, in this experiment, the advantages of structural prediction algorithms are not fully exploited, making the setup more favorable for the algorithms in previous studies. It is also worth noting that by using the same decoding function as theirs, our approach can achieve the same order of the cumulative losses in online multiclass classification with bandit feedback.

### v2

@@ H.2 Multilabel classification
As the algorithm $\mathsf{ALG}$ to update the linear estimator, we employ the OGD in Section 3.2 with learning rate $\eta_{t}=\frac{B}{\sqrt{2\left(10^{-8}+\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}\right)}}$ and orthogonal projection onto a ball of diameter 50. Here, the addition of $10^{-8}$ to the denominator is to prevent division by zero. We fixed $B=50$ for the experiments. All other parameters were set according to theoretical values. Under these parameter settings, we repeated experiments $10$ times.
Results
We provide the results in Figure 3. When $d$ is small, the loss incurred by the algorithm based on the inverse-weighted gradient estimator is relatively low, whereas when $d$ is large, the algorithm based on the pseudo-inverse matrix estimator yields a smaller loss. The superiority of the algorithm based on the inverse-weighted gradient estimator for small $d$ is consistent with the theoretical result that this algorithm has a better dependence on $T$. Similarly, the better performance of the algorithm based on the pseudo-inverse matrix estimator for large $d$ matches the theoretical result that this algorithm does not depend explicitly on $K$. Thus, these experimental findings provide empirical validation of our theoretical results.